# Majorana vortex phases in time-reversal invariant higher-order topological insulators and topologically trivial insulators

### v2

@@ VII Acknowledgments
Xun-Jiang Luo thanks K. T. Law, Fengcheng Wu, and Zhongyi Zhang for helpful discussions.
@@ Appendix A Surface states projection analysis
For $31<m_{0}<$ 3, the term $\tau_{0}\otimes H_{\text{TI}}$ exhibits a double band inversions at the $\Gamma$ point, resulting in two gapless surface Dirac cones. These surface states can be gapped by additional mass terms. For simplicity, we set $\lambda=0$ in the following analysis and consider the (100) surface as an example. By expanding the Bloch Hamiltonian around the $\Gamma$ point and imposing open boundary conditions along the $x$-direction, we obtain $H_{\mathrm{HOTI}}^{(1)}=h_{0}+h_{p}$, where:
@@
with $\tilde{m}=m_{0}-3$ and $\tilde{\eta}=\eta_{0}+\eta_{1}$. The Hamiltonian $h_{0}$ hosts eight zero-energy states, which are eigenstates of the operator $\tau_{0}\sigma_{y}s_{x}$. The wavefunctions of these zero-energy states can be analytically derived by solving the eigenvalue equation $h_{0}\psi_{\alpha}(x)=0$ [39]. By projecting the perturbation Hamiltonian $h_{p}$ onto the subspace spanned by the zero-energy states localized at $x=L$, we obtain the Dirac mass for the surface states on the (100) surface, given by $\Delta_{x}=\eta_{0}+\tilde{m}\eta_{1}$. Similarly, the Dirac masses for the (010) and (001) surfaces are $\Delta_{y}=\eta_{0}-\tilde{m}\eta_{1}$ and $\Delta_{z}=\eta_{0}$, respectively. For $m_{0}=2.5$, these become $\Delta_{x}=\eta_{0}-\eta_{1}/2$, $\Delta_{y}=\eta_{0}+\eta_{1}/2$, and $\Delta_{z}=\eta_{0}$. Thus, for $20<\eta_{0}<\eta_{1}/$ 2, the (100) surface develops a gap with a sign opposite to that of the (010) and (001) surfaces, leading to HHMs, as illustrated in Fig. 1(c).
@@ Appendix B MZMs protected by rotational symmetry
[…]
@@ Appendix D Effective model Hamiltonian of HOTI Bi
[…]

### v3

@@ VII Acknowledgments
Xun-Jiang Luo thanks Fengcheng Wu, K. T. Law, and Zhongyi Zhang for helpful discussions.
@@ Appendix A Surface states projection analysis
For $1<m_{0}<3$, the term $\tau_{0}\otimes H_{\text{TI}}$ exhibits a double band inversions at the $\Gamma$ point, resulting in two gapless surface Dirac cones. These surface states can be gapped by additional mass terms. For simplicity, we set $\lambda=0$ in the following analysis and consider the (100) surface as an example. By expanding the Bloch Hamiltonian around the $\Gamma$ point and imposing open boundary conditions along the $x$-direction, we obtain $H_{\mathrm{HOTI}}^{(1)}=h_{0}+h_{p}$, where:
@@
with $\tilde{m}=m_{0}-3$ and $\tilde{\eta}=\eta_{0}+\eta_{1}$. The Hamiltonian $h_{0}$ hosts eight zero-energy states, which are eigenstates of the operator $\tau_{0}\sigma_{y}s_{x}$. The wavefunctions of these zero-energy states can be analytically derived by solving the eigenvalue equation $h_{0}\psi_{\alpha}(x)=0$ [88]. By projecting the perturbation Hamiltonian $h_{p}$ onto the subspace spanned by the zero-energy states localized at $x=L$, we obtain the Dirac mass for the surface states on the (100) surface, given by $\Delta_{x}=\eta_{0}+\tilde{m}\eta_{1}$. Similarly, the Dirac masses for the (010) and (001) surfaces are $\Delta_{y}=\eta_{0}-\tilde{m}\eta_{1}$ and $\Delta_{z}=\eta_{0}$, respectively. For $m_{0}=2.5$, these become $\Delta_{x}=\eta_{0}-\eta_{1}/2$, $\Delta_{y}=\eta_{0}+\eta_{1}/2$, and $\Delta_{z}=\eta_{0}$. Thus, for $0<\eta_{0}<\eta_{1}/2$, the (100) surface develops a gap with a sign opposite to that of the (010) and (001) surfaces, leading to HHMs, as illustrated in Fig. 1(c).
@@ Appendix B MZMs protected by rotational symmetry
$\gamma=(e^{-i\pi/4}c_{\downarrow}+e^{i\pi/4}c_{\downarrow}^{\dagger})e^{-\int_{0}^{r}\Delta(r^{\prime})/v_{F}dr^{\prime}},$ (11)
@@ Appendix D Effective model Hamiltonian of HOTI Bi
Here, $\mu\in\{1,\ \cdots,5\}$, $i,j\in\{\mathrm{I},\mathrm{II}\}$, $\nu\in\{1,\ 2\}$, $C^{z}_{3,\mathrm{I}}=\sigma_{0}\otimes e^{\mathrm{i}\frac{\pi}{3}\sigma_{3}}$ and $C^{z}_{3,\mathrm{II}}=-\sigma_{0}\otimes\sigma_{0}$ so that the full threefold rotation symmetry is given by $C_{3z}=C^{z}_{3,\mathrm{I}}\oplus C^{z}_{3,\mathrm{II}}$. The lattice basis is $\bm{a}_{1}=(1,0,0)$, $\bm{a}_{2}=(-1/2,\sqrt{3}/2,0)$, and $\bm{a}_{3}=(0,0,1)$. Following Ref. [69], we take $m_{\mathrm{I}}=m_{\mathrm{II}}=2,t_{\mathrm{I}}=t_{\mathrm{II}}=1,\lambda_{\mathrm{I}}=0.3,\lambda_{\mathrm{II}}=\gamma_{\mathrm{II}}=1,\epsilon=0.1\text{, and }\delta=0.3$.